A regional airline is the only carrier on a local air route and must determine the number of flights it will provide per week, and the fare it will charge. The estimated cost - fuel, airport charges, pilots, etc. - per flight is $2,000, and each flight is expected to be full, with 100 passengers. Thus the marginal cost per passenger may be considered to be $20.The estimated demand curve is P = 120 − 0.1Q, where P is the fare in dollars and Q is the number of passengers per week.

(a) What is the airline’s profit-maximizing fare?

(b) How many passengers does it carry per week, and in how many flights?

(c) What is the weekly profit?

(d) The airline is offered $4,000 per week to haul freight along the route. However, this will mean replacing one of the weekly passenger flights with a freight flight (at the same operating cost). Should the airline carry freight for the local firm? Explain
Answer #1

P=120-0.1Q

MR=120-0.2Q

Profit maximising condition MC=MR

120-0.2Q=20

Q*=100/0.2=500

P*=120-0.1*500=$70

Proft=TR-TC= P*Q-20*Q= 500*70-20*500=35000-10000=$25000

Answer-

a. Profit maximising fare=$70

b. b. Airline carries 500 passenger per week and 5 flights.

c. Weekly profit=$25000

d. No, It should not accept the offer as the profit for 1 flight is $25000/5=$5000 and the airline is offered $4000 only. So accepting the offer means the airline suffers a loss of $5000-$4000=$1000. Hence the airline should not accept the offer.